# A First Look at Bugs in LLM Inference Engines

### v1

@@ 4. RQ1: Symptoms
To clarify the impact of bugs and provide insights for test oracle, we explore the bug symptoms and their distribution of LLM inference engines.
@@ 4.1. Symptom Classification
As shown in Table 2, we classify the observed symptoms into six unique categories, including Crash (S1), Unexpected Output (S2), Feature Failure (S3), Abnormal Performance (S4), System Hang (S5), and Silent Error (S6). These symptoms vary in distribution across three phases of LLM deployment workflow (i.e., engine setup, model conversion, and inference/serving), as shown in Table 3. Next, we illustrate the symptoms by phase.
{mdframed}
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Finding 1: Six bug symptoms are identified, including crash, unexpected output, feature failure, abnormal performance, system hang, and silent error.
@@ 4.1.1. Engine Setup
{mdframed}
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Finding 2: Most (97%) bugs in engine setup phase lead to a crash.
@@
S2 Unexpected Output Engines produce unexpected or incorrect outputs.
@@
S4 Abnormal Performance Engines significantly deviate from expected performance (e.g., resource usage and response latency).
@@ 4.1.2. Model Conversion
{mdframed}
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Finding 3: In model conversion phase, 83% of bugs manifest as failures to generate model files (S1, S5) while 17% produce incorrect files (S3, S6).
@@ 4.1.3. Inference/Serving
We observe 664 bugs (71% of total) in the inference/serving phase, covering all six identified symptoms.
Most (59%) bugs cause the engine to stop functioning, involving symptoms of S1 and S5. Crash (S1) is the most common symptom, with 363 instances involving initialization failures or irrecoverable service termination. For example, in vLLM issue #8952, the server crashed when receiving requests with configurations like ‘response_format’ or ‘guided_json’ System Hang (S5) occurs in 30 cases, where the service becomes unresponsive without crashing, as seen in Mlc-llm issue #306, where API calls hang during prefill function execution.
93 bugs make engines exhibit functional anomalies while remaining operational, which are classified into Feature Failure (S3). Common issues include abnormal API response formats or task execution failures with user-specific configurations. For instance, in vLLM’s issue #8947, the engine mishandles the configuration priority logic.
Bugs can also happen even if engines generate output tokens, exhibiting as S2, S4, and S5. Unexpected Output (S2), with 121 instances, occurs when the service produces user-unexpected tokens or metric values (e.g., probability distribution). For example, in the DeepSpeed issue #3285, the tensor computation results of the engine differ from those of the original PyTorch model. showing parameter size consistency with PyTorch but logits deviations Abnormal Performance (S4), with 53 bugs, manifests as either excessively high or low performance metrics (e.g., latency, throughput, and resource utilization). For instance, issue #2488 of DeepSpeed reports an unexpected increased text generation latency. Silent Error (S6), with 4 bugs, involves hidden code defects that may disrupt computations. For example, the issue #6030 of vLLM reports that an if-else condition branches contained identical code blocks.
{mdframed}
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Finding 4: The inference/serving phase has the highest bug proportion (71%) and exhibits all six symptom types.
@@
Following the labeling process in section 3, among the 121 ‘Unexpected Output’ bugs in inference/serving phase, we further identify 102 bugs [^2] The other 19 bugs are not associated with output tokens, such as DeepSpeed issue #3285. with unexpected output tokens, covering 5 factors, whose distribution is shown in Fig. 4. The most (36%) observed factor is unexpected Character (UT1), with 37 bugs, which indicates that the output tokens of LLMs involve massive garbled text (e.g. Llama.cpp issue #1735). Semantic Coherence (UT2) is the second most (20%) used factor, found in 20 cases, implying that the outputs of LLM are meaningless (e.g. Llama.cpp issue #3782). Inconsistent Output (UT3) indicates that even after disabling all randomness settings, the LLM outputs still vary across different optimization algorithms or runtime environments. This factor accounts for 17%, with 17 instances. For instance, in vLLM issue #590, output tokens differ between the original model and the model served using vLLM. 15% of bugs, found in 15 issues, are identified by unexpected Length (UT4) of output tokens, occur when outputs are unusually short or excessively long, never stopping, exemplified by vLLM issue #9448. Repetition (UT5) issues, seen in 13 cases, involve excessive duplication of strings. For example, in DeepSpeed issue #4903, the generated tokens abnormally repeat the ‘e’ character.
{mdframed}
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Finding 5: Despite non-deterministic outputs of LLMs, five factors (i.e. character, semantic coherence, inconsistent output, length, and repetition) provide insights for oracle settings.
4.2. Symptom Distribution
Overall, Crash (S1) is the most prevalent (65%) symptom across all three phases. Unexpected Output (S2) ranks second at 13%, primarily occurring during the inference/serving phase. Feature Failure (S3) follows at 11%, with Performance Anomalies at 6%, System Hangs at 4%, and Silent Errors at 1%. Among these symptoms, crashes are commonly used for bug detection in DL libraries (Deng et al., 2023, 2024) due to their distinct characteristics and easy detection through core dumps or stack traces. Moreover, unexpected outputs are the second most common symptoms, comprising nearly 20% during the inference/serving phase. This underscores the need for suitable test oracles for LLM outputs and highlights the significance of Finding 4.
{mdframed}
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Finding 6: While crashes are the most common, over 35% of bugs show other symptoms, necessitating diverse oracles beyond crashes for LLM inference engine testing.
@@ 5.1. Root Cause Classification
{mdframed}
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Finding 7: We develop a taxonomy of 28 specific root causes of bugs in LLM inference engines, indicating the diversity of bug sources in these complex engines.
@@
For user input, inference engines must parse the formatted job configurations and prompts. The Incorrect Request Parsing (RA.2), with 10 cases, indicates a failure to validate and parse requests, or provide feedback for malformed inputs. For example, in Llama.cpp issue #3766, a request with malformed ‘system_prompt’ configuration caused a system hang without warning due to inadequate request validation. Additionally, in this process, engines must avoid Incorrect Character Encoding (RA.6), with 2 cases, to accurately encode the characters of prompts. RA.6 highlights the mishandling of incompatibilities between the engine’s and users’ character encodings. For instance, in Mlc-llm issue #804, the engine struggled with parsing some Unicode/UTF-8 characters of user prompts correctly.
After parsing requests, engines will tokenize the prompt for subsequent inference. Incorrect Tokenization (RA.1), with 12 cases, indicates the use of misconfigured tokenizer or incorrect tokenization algorithm, leading to wrong token splits. For instance, in Llama.cpp issue #290, the tokenizer improperly tokenized strings like ‘china’. RA.1 is the primary cause of in-/output relevant bugs, underscoring the need for meticulous attention to tokenizer checks, handling, and error tolerance. Moreover, engines face challenges in parsing multimodal data, often encountering issues with format parsing and dimension alignment for multimodal inputs, i.e., Insufficient Multimodal Support (RA.3) with 9 cases. For example, in vLLM issue #9128, the engine failed to correctly merge image and text embeddings.
@@
{mdframed}
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Finding 8: Although in-/output relevant bugs are less frequent (comprising 4% of bugs), they significantly impact user interaction, necessitating careful checking, processing, and error handling of character encoding, tokenizer, and user-specific formatting.
@@
Moreover, to improve resource utilization and accelerate inference, engines are optimized for handling multiple requests under diverse concurrency configurations, which is also the major source of bugs, with 42 cases. Concurrency Issues (RB.5), with 39 cases, implies the incorrect management of concurrent workers. For instance, in Llama.cpp issue #3537, the static variable ‘mu’ used by the ‘Mirostat’ sampler was shared across different workers, leading to state conflicts. This implies the need for engines to carefully address potential concurrency issues, including shared variables, atomic operations, and resource contention. Moreover, Incorrect Job Context Management (RB.6), with 3 cases, refers to errors in handling context and configurations for multiple requests. For instance, in TensorRT-LLM issue #1447, the engine fails to process requests with and without the ‘SamplingConfig’ parameter simultaneously.
Moreover, Document Issues (RB.5), with 13 cases, refer to inaccurate or outdated descriptions of usage and configurations in documentation. For example, DeepSpeed issue #4039 highlights the absence of inference configuration details for MoE models.
{mdframed}
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Finding 9: Managing configurations is challenging for LLM inference engines, accounting for 23% of bugs across six root causes. Engines should prioritize comprehensive compatibility testing and timely updates for configuration combinations, model adaptation, mixed precision, concurrency conditions, multitasking, and documentation.
@@
Bug can originate from root causes that are common in DL libraries. Incorrect Algorithm Implementation (RC.1) refers to problematic functional implementation, such as the optimization strategies and computing kernels. For instance, in vLLM issue #590, an erroneous implementation of ‘RoPE’ algorithm within attention operators caused unexpected output tokens. This represents the most (43%) prevalent functionality issue with 184 cases, emphasizing the complexity of functionalities in LLM inference engines. Misused API (RC.2), with 75 cases, refers to incorrect usage of either internal or external APIs. This includes (1) failing to use an API or using it redundantly, and (2) using an incorrect API name or arguments. For example, in vLLM issue #9059, directly accessing a non-existent ’vocab’ attribute in ’MistralTokenizer’ instead of using ’get_vocab()’ caused a bug. Mismatched Shape (RC.3), with 53 cases, occurs when tensor dimensions violate operator requirements. For instance, in TensorRT-LLM issue #2150, a ‘5D’ tensor input to conv2d layers (which expect ‘3D/4D’ inputs) led to a crash.
@@
{mdframed}
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Finding 10: The complex functionality of LLM inference engines contributes to 41% of bug root causes. The diversity of nine failure modes highlights the need for formal specification checking, comprehensive unit testing, and runtime validation across components of engines.
@@
{mdframed}
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Finding 11: Environmental compatibility poses significant challenges for LLM inference engines, accounting for 29% cases, spanning three categories. This underscores the need for thorough compatibility testing, efficient version management, and reliable dependency management across various backends.
@@
{mdframed}
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Finding 12: Resource management bugs, though only 4%, are critical due to risks like memory leaks. We identify four leaf root causes, highlighting the need for carefully checking on resource allocation and release, cache management, and multi-device management.

### v2

@@ 3.3.2. Reliability Analysis
3.3.3. Fix Strategy Analysis
Based on the constructed root cause taxonomy, we extend our study to include bug fix strategies for the root causes. The methodology follows the same rigorous procedure as the symptom and root cause labeling.
Utilizing the pilot labeling dataset mentioned earlier, we employ the GitHub API (git, 2025) to retrieve the associated fix commits for bugs linked to each internal root cause category (see Section 5). Then, the two inspectors independently read the contents of the retrieved fix commits, including the title, developer conversations, and changed files. Next, they assign initial fix strategies to each fix commit based on the contents. They then engage in discussions to reach agreements on the initial codes. The arbitrator is involved to address any disagreement. Subsequently, the three authors iteratively refine the initial codes by grouping similar codes and continuously reviewing the categories and contents of the fix commits. Finally, they obtain initial fix strategies for bugs associated with each internal root cause category.
For reliability analysis, following the same procedure, two inspectors label the fix commits of remaining issues based on the initial codes of fix strategies. The $\kappa$ value for the fix strategy labeling is $0.8179$, indicating almost perfect agreement. No new category of fix strategies is introduced. The categories that fix at least 5 issues are identified as common fix strategies.
@@ 4. RQ1: Symptoms
To clarify the impact of bugs and provide insights for test oracle, we explore the bug symptoms and their distribution across LLM inference engines.
@@ 4.1. Symptom Classification
We classify the observed symptoms into six unique categories, as shown in Table 2. Crashes (S1) refer to unexpected engine termination without recovery capability, such as the unrecoverable termination of compilation or inference process due to out-of-bound memory access (Llama.cpp issue #10176). Unexpected Output (S2) denotes the generation of incorrect model parameters, garbled tokens, or responses that are illogical or irrelevant to the prompt. An example of this is the generation of a random mix of words in different languages, as reported in Llama.cpp issue #1735. Feature Failure (S3) refers to the engine remaining running but failing to perform specific features, such as the invalid aliases configured for the model instance noted in vLLM issue #8947. Abnormal Performance (S4) refers to token generation with unusually high or low resource consumption or latency, like the absence of GPU consumption and increased latency in Llama.cpp issue #3806. System Hang (S5) indicates that the engine becomes unresponsive but does not terminate, such as when the model quantization process stalls without completion, as seen in TensorRT-llm issue #1356. Silent Error (S6) involves internal anomalies that evade detection mechanisms. For example, in the Llama.cpp issue #9245, the engine produced corrupted model files without generating any warnings.
Finding 1: Six bug symptoms are identified, including crash, unexpected output, feature failure, abnormal performance, system hang, and silent error.
4.2. Symptom Distribution
We study the distribution of identified symptoms across the three phases of LLM inference engine deployment workflow (i.e., engine setup, model conversion, and inference/serving). As shown in Table 3, these symptoms vary in distribution across these phases. Note that although different phases may exhibit the same symptom (e.g., a crash), their impacts could differ. For instance, a crash during the setup phase may prevent an engine from being installed in the specific environment. Meanwhile, a crash during model conversion results in a faulty model, rendering it unusable during the inference/serving stage. Additionally, a crash during inference/serving stage disrupts user interaction and service availability.
@@ 4.2.1. Engine Setup
Finding 2: Most (97%) bugs in engine setup phase lead to a crash.
@@
S2 Unexpected Output Engines output incorrect model parameters or tokens that are unreadable, incoherent, unrelated to prompts, or fail to maintain logical capability.
@@
S4 Abnormal Performance Engines generate tokens with excessive or insufficient resource usage, or exhibit abnormal generation latency.
@@ 4.2.2. Model Conversion
Finding 3: In model conversion phase, 83% of bugs prevent the generation of model files (S1, S5) while 17% lead to the creation of incorrect model files (S3, S6).
@@ 4.2.3. Inference/Serving
We observe 664 bugs (71% of total) in the inference/serving phase, covering all six identified symptoms. We describe them in a gradual order, beginning with the most disruptive failures, moving to cases where the engine continues to run but behaves incorrectly, and concluding with other symptoms observed when outputs are still produced.
Service-disrupting bugs. Most (59%) bugs cause the engine to stop functioning, involving symptoms of S1 and S5. Crash (S1) is the most common symptom, with 363 instances involving initialization failures or irrecoverable service termination. For example, in vLLM issue #8952, the server crashed when receiving requests with configurations like ‘response_format’ or ‘guided_json’. System Hang (S5) occurs in 30 cases, where the service becomes unresponsive without crashing, as seen in Mlc-llm issue #306, where API calls hang during prefill function execution.
Functional anomalies bugs. 93 bugs make engines exhibit functional anomalies while remaining operational, which are classified into Feature Failure (S3). For instance, in vLLM issue #8947, the engine’s argument parser incorrectly required the alias (‘served-model-name’) to be the first parameter. A configuration where the alias appeared later, such as ‘(host=‘xxx’, port: ‘xxx’, served-model-name=‘MyModel’)’, would cause the engine to erroneously ignore the alias (served-model-name=‘MyModel’) silently. In this case, vLLM performed model inference without setting the configured alias. Therefore, this was classified as a Feature Failure.
Bugs during output generation. Bugs can also happen when engines start generating output tokens, exhibiting as S2, S4, and S5. Unexpected Output (S2), with 121 instances, occurs when the service produces user-unexpected tokens or metric values (e.g., probability distribution). This category consists of two subtypes: (i) unexpected tokens from LLM generation and (ii) unexpected non-token outputs. Unexpected tokens refer to cases where the LLM-generated tokens are not human-readable, semantically coherent, logically consistent, or prompt-relevant, such as producing gibberish or irrelevant text. For example, Llama.cpp issue #1735 reported unreadable results (e.g., ‘ &amp Vertigowebsitesearch engines Search engines Search’), which were classified as unexpected tokens. The second subtype includes non-token outputs, which occur when components like setup tools, model conversion tools, and monitoring tools within the engines generate erroneous setup files, model files, model parameter values, or metric values. For example, Llama.cpp issue #10285 reported a case of incorrect output of model parameter and size counts due to incorrect algorithm implementation. Showing parameter size consistency with PyTorch but logits deviations Abnormal Performance (S4), with 53 bugs, manifests as either excessively high or low performance metrics (e.g., latency, throughput, and resource utilization). For instance, issue #2488 of DeepSpeed reports a case of abnormal performance, where the supposedly optimized DeepSpeed model was slower than the vanilla model, with its average latency increasing from ~8566ms to ~8811ms. Silent Error (S6), with 4 bugs, involves hidden code defects that may disrupt computations. For example, the issue #6030 of vLLM reports that an if-else condition branch contained identical code blocks.
Finding 4: The inference/serving phase has the highest bug proportion (71%) and exhibits all six symptom types.
@@
Following the labeling process in section 3, among the 121 ‘Unexpected Output’ bugs in inference/serving phase, we further identify 102 bugs  [^2] The remaining 19 cases involve unexpected non-token outputs. with unexpected output tokens, covering 5 factors, whose distribution is shown in Fig. 4. Among the unexpected token output, the most (36%) observed factor is unexpected Character (UT1), with 37 bugs, which indicates that the output tokens of LLMs involve massive garbled text (e.g. Llama.cpp issue #1735). Semantic Coherence (UT2) is the second most (20%) used factor, found in 20 cases, implying that the outputs of LLM are meaningless (e.g. Llama.cpp issue #3782). Inconsistent Output (UT3) indicates that even after disabling all randomness settings, the LLM outputs still vary across different optimization algorithms or runtime environments. This factor accounts for 17%, with 17 instances. For instance, in vLLM issue #590, output tokens differ between the original model and the model served using vLLM. 15% of bugs, found in 15 issues, are identified by unexpected Length (UT4) of output tokens, occur when outputs are unusually short or excessively long, never stopping, exemplified by vLLM issue #9448. Repetition (UT5) issues, seen in 13 cases, involve excessive duplication of strings. For example, in DeepSpeed issue #4903, the generated tokens abnormally repeat the ‘e’ character.
Finding 5: Despite non-deterministic outputs of LLMs, five factors (i.e. character, semantic coherence, inconsistent output, length, and repetition) provide insights for oracle settings.
Overall, Crash (S1) is the most prevalent (65%) symptom across all three phases. Unexpected Output (S2) ranks second at 13%, primarily occurring during the inference/serving phase. Feature Failure (S3) follows at 11%, with Abnormal Performance (S4) at 6%, System Hang (S5) at 4%, and Silent Error (S6) at 1%. Among these symptoms, crashes are commonly used for bug detection in DL libraries (Deng et al., 2023, 2024) due to their distinct characteristics and easy detection through core dumps or stack traces. Moreover, unexpected outputs are the second most common symptoms, comprising nearly 20% during the inference/serving phase. This underscores the need for suitable test oracles for LLM outputs and highlights the significance of Finding 4.2.3.
Finding 6: While crashes are the most common, over 35% of bugs show other symptoms, necessitating diverse oracles beyond crashes for LLM inference engine testing.
@@ 5.1. Root Cause Classification
Finding 7: We develop a taxonomy of 28 specific root causes of bugs in LLM inference engines, indicating the diversity of bug sources in these complex engines.
@@
For user input, inference engines must parse the formatted job configurations and prompts. The Incorrect Request Parsing (RA.2), with 10 cases, indicates a failure to validate and parse requests, or provide feedback for malformed inputs. For example, in Llama.cpp issue #3766, a request with malformed ‘system_prompt’ configuration caused a system hang without warning due to inadequate request validation. Additionally, in this process, engines must avoid Incorrect Character Encoding (RA.6), with 2 cases, to accurately encode the characters of prompts. RA.6 highlights the mishandling of incompatibilities between the engine’s and users’ character encodings. For instance, in Mlc-llm issue #804, the engine failed to correctly output certain characters (such as emoji and CJK characters).
After parsing requests, engines will tokenize the prompt for subsequent inference. Incorrect Tokenization (RA.1), with 12 cases, indicates the use of misconfigured tokenizer or incorrect tokenization algorithm, leading to wrong token splits. For instance, in Llama.cpp issue #290, the tokenizer completely failed to process the input string ‘ china’ (a space ‘ ’ and a word ‘china’). It reported only one token corresponding to the space (‘ ’) in the prompt while ignoring the actual word (‘china’), leading to an empty effective input for the model. RA.1 is the primary cause of in-/output relevant bugs, underscoring the need for meticulous attention to tokenizer checks, handling, and error tolerance. Moreover, engines face challenges in parsing multimodal data, often encountering issues with format parsing and dimension alignment for multimodal inputs, i.e., Insufficient Multimodal Support (RA.3) with 9 cases. For example, in vLLM issue #9128, the engine failed to correctly merge image and text embeddings.
@@
Finding 8: Although in-/output relevant bugs are less frequent (comprising 4% of bugs), they significantly impact user interaction, necessitating careful checking, processing, and error handling of character encoding, tokenizer, and user-specific formatting.
@@
Moreover, to improve resource utilization and accelerate inference, engines are optimized for handling multiple requests under diverse concurrency configurations, which is also the major source of bugs, with 42 cases. Concurrency Issue (RB.3), with 39 cases, implies the incorrect management of concurrent workers. For instance, in Llama.cpp issue #3537, the static variable ‘mu’ used by the ‘Mirostat’ sampler was shared across different workers, leading to state conflicts. This implies the need for engines to carefully address potential concurrency issues, including shared variables, atomic operations, and resource contention. Moreover, Incorrect Job Context Management (RB.6), with 3 cases, refers to errors in handling context and configurations for multiple requests. For instance, in TensorRT-LLM issue #1447, the engine fails to process requests with and without the ‘SamplingConfig’ parameter simultaneously.
Moreover, Document Issue (RB.5), with 13 cases, refers to inaccurate or outdated descriptions of usage and configurations in documentation. For example, DeepSpeed issue #4039 highlights the absence of inference configuration details for MoE models.
Finding 9: Managing configurations is challenging for LLM inference engines, accounting for 23% of bugs across six root causes. Engines should prioritize comprehensive compatibility testing and timely updates for configuration combinations, model adaptation, mixed precision, concurrency conditions, multitasking, and documentation.
@@
Bug can originate from root causes that are common in DL libraries. Incorrect Algorithm Implementation (RC.1) refers to problematic functional implementation, such as the optimization strategies and computing kernels. For instance, in vLLM issue #590, an erroneous implementation of ‘RoPE’ algorithm within attention operators caused unexpected output tokens. This represents the most (43%) prevalent functionality issue with 184 cases, emphasizing the complexity of functionalities in LLM inference engines. Misused API (RC.2), with 75 cases, refers to incorrect usage of either internal or external APIs. This includes (1) failing to use an API or using it redundantly, and (2) using an incorrect API name or arguments. For example, in vLLM issue #9059, directly accessing a non-existent ’vocab’ attribute in ’MistralTokenizer’ instead of using ’get_vocab()’ caused a bug. Additionally, RC.1 is used when the core problem lies in the algorithmic logic (e.g., wrong formula or algorithm adoption), while RC.2 is used when the bug stems from incorrect API usage (e.g., wrong arguments, wrong function calls). Mismatched Shape (RC.3), with 53 cases, occurs when tensor dimensions violate operator requirements. For instance, in TensorRT-LLM issue #2150, a ‘5D’ tensor input to conv2d layers (which expect ‘3D/4D’ inputs) led to a crash.
@@
Finding 10: The complex functionality of LLM inference engines contributes to 41% of bug root causes. The diversity of nine failure modes highlights the need for formal specification checking, comprehensive unit testing, and runtime validation across components of engines.
@@
Finding 11: Environmental compatibility poses significant challenges for LLM inference engines, accounting for 29% cases, spanning three categories. This underscores the need for thorough compatibility testing, efficient version management, and reliable dependency management across various backends.
@@
Finding 12: Resource management bugs, though only 4%, are critical due to risks like memory leaks. We identify four leaf root causes, highlighting the need for carefully checking on resource allocation and release, cache management, and multi-device management.
The substantial memory and computational requirements of LLMs lead to a pronounced prevalence of bugs stemming from two main root causes. The first category comprises bugs directly resulting from resource management mechanisms—such as KV-cache allocation or memory pool scheduling—which we classify as ‘Resource’ (RE). The second category is ‘Functionality’ (RC), which involves bugs where incorrect functional implementation induces resource-related problems, such as memory leaks due to flawed kernel logic. Empirical data suggests that bugs caused purely by resource management mechanisms are relatively infrequent, as RE accounts for only 4% of observed cases. In contrast, the majority of resource-related bugs arise from faulty functional implementations, as functionality-related issues dominate our taxonomy, comprising 41% of all bugs. These findings indicate that the notorious memory demands of LLMs significantly increase the complexity and vulnerability of functional implementations and resource management.
Table 4. Distribution of root cause types and the most frequent root causes across deployment phases. The numbers in parentheses indicate the number of issues and their respective percentages.
Phase # of Root Top1 Top2 Top3 Top4 Top5
Cause Types
Engine Setup 12 (RD.1) Incompa- tible Backend (40, 26.0%) (RD.3) Dependent Module Issue (35, 22.7%) (RB.1) Misconfigu- ration (30, 19.5%) (RD.2) Incompa- tible Version (22, 14.3%) (RC.2) Misused API (6, 3.9%)
Model Conversion 15 (RC.1) Incorrect Algorithm Implementation (23, 14.6%) (RB.2) Incompa- tible Model (18, 11.4%) (RB.1) Misconfigu- ration (15, 9.5%) (RD.1) Incompa- tible Backend (13, 8.2%) (RC.3) Mismat -ched Shape (13, 8.2%)
Inference/ Serving 28 (RC.1) Incorrect Algorithm Implementation (158, 21.3%) (RD.1) Incompa- tible Backend (85, 11.5%) (RD.2) Incompa- tible Version (62, 8.4%) (RB.1) Misconfigu- ration (60, 8.1%) (RC.2) Misused API (56, 7.6%)
5.2. Root Cause Distribution
Table 4 summarizes the distribution of root causes across different deployment phases.
Bugs in the Engine Setup phase stem from a total of 12 distinct types of root causes. The primary root causes include (RD.1) Incompatible Backend, (RD.3) Dependent Module Issue, (RB.1) Misconfiguration, (RD.2) Incompatible Version, and (RC.2) Misused API. This indicates that the engine setup phase requires focused attention on ensuring environment compatibility and rigorous configuration management.
Bugs in the Engine Setup phase are caused by 15 unique root cause types. The main root causes of bugs are (RC.1) Incorrect Algorithm Implementation, (RB.2) Incompatible Model, (RB.1) Misconfiguration, (RD.1) Incompatible Backend, and (RC.3) Mismatched Shape. This suggests that model conversion requires particularly careful algorithm implementation, configuration management, and environment adaptation.
Bugs in the Inference/Serving phase involve the highest number of root cause types, totaling 28. The predominant root causes are (RC.1) Incorrect Algorithm Implementation, (RD.1) Incompatible Backend, (RD.2) Incompatible Version, (RB.1) Misconfiguration, and (RC.2) Misused API. Similar to the model conversion phase, algorithm implementation is also a major challenge in the inference/serving phase, followed by environment adaptation and configuration management.